How can I Find the solutions to the equation: x− 5/x = 4

Answer by stanbon(75887)

Find the solutions to the equation: x− 5/x = 4
Multiply thru by x to get:
x^2-5 = 4x
Rearrange
x^2-4x-5=0
Factor
(x-5)(x+1)=0
x=5 or x=-1
================
Cheers,
Stan H.